# Hierarchical Learning and Computing over Space-Ground Integrated Networks

### v1

@@ III-B Problem Formulation
where $w_{i,j}(m)$ can be regarded as the weight of edge $(i,j)$ in $\mathcal{E}^{\prime}(m)$. This is equivalent to finding a minimum directed sub-branching covering all the terminals with the lowest energy consumption for a given topology. Essentially, it involves identifying an inter-satellite routing path at minimum cost that takes the form of a directed sub-branching.
@@
to minimize the energy consumption in each time frame. Here $\mathcal{B}(m,u)$ denotes the directed sub-branching of each time frame. This means that the solution to $\mathscr{P}_{1}$ can be obtained by calculating the solution to each time frame by solving $\mathscr{P}_{2}$. It is worth noting that the local information on each satellite is assumed to be partitioned into $U$ parts, so that each part can be transmitted within the duration of a time frame. In this case, the energy consumption for ISLs in each time frame is given by
@@
The overall characterization of space topology is demonstrated in Fig. 3.
@@ III-C Proposed Topology-Aware Energy-Efficient Routing
    Shortest Path: Find the shortest distance path from each terminal node $i\in\mathcal{V}^{\prime}\setminus r$ to the root node $j=r$ using the Dijkstra algorithm [27], and record each predecessor node such that each terminal can trace back to the root node. Construct a set $\mathrm{Path}=\{i\rightarrow r|i\in\mathcal{V}\setminus r\}$ consisting of node and edge information for all shortest paths.
@@
The detailed procedures of the proposed TAEER algorithm are presented in Algorithm 1 and an simple example of the solution is illustrated in Fig. 4. Based on the derived MSA $\mathcal{A}^{\ast}(m,u)$, a tree structure in such an LEO satellite network enables transmission-and-aggregation from each leaf node to the root node, ultimately completing the global model aggregation task.
@@ III-D Extension to Unpredictable Topology
We consider transmission failure caused by outage in the laser ISL. To characterize the outage probability of the laser ISL, it can be modeled as the probability that the signal-to-noise ratio (SNR) falls below a certain threshold, which can be expressed as
@@
To minimize the overall outage probability of the aggregation path, the optimization problem can be derived as
$\mathscr{P}_{3}:~{}\operatorname{minimize}~{}$ $P_{\text{out}}(\mathcal{B}(m,u))=\prod_{i,j\in\mathcal{V}^{\prime}}p_{\text{out}}^{i,j}(m,u).$ (36)
@@
$\mathscr{P}_{4}:~{}\operatorname{minimize}~{}$ $\log P_{\text{out}}(\mathcal{B}(m,u))=\sum_{i,j\in\mathcal{V}^{\prime}}\log p_{\text{out}}^{i,j}(m,u).$ (37)
As a result, to balance the energy efficiency and the outage probability of the system, the objective function can be formulated as the linear combination of $\mathscr{P}_{2}$ and $\mathscr{P}_{4}$, i.e.,
$\mathscr{P}_{5}:~{}\operatorname{minimize}~{}$ $\rho W_{u}(\mathcal{B}(m,u))+(1-\rho)\log P_{\text{out}}(\mathcal{B}(m,u)).$ (38)
It is worth noting $\mathscr{P}_{5}$ can be directly solved by adjusting the weight of the directed graph to $w^{r}_{i,j}=\rho w_{i,j}+(1-\rho)\log P_{\text{out}}^{i,j}$ using Algorithm 1.
@@ IV-A1 Network Configurations
To verify the performance of the proposed algorithms, we consider a $80/4/1$ Walker-Star constellation system and a $80/4/1$ Walker-Delta constellation system, which consists of $N=4$ orbit planes and $S_{n}=20$ satellites evenly distributed in each orbit for default settings. Three GEO satellites are evenly distributed in the Equatorial plane orbit. In order to characterize the impact of the network size, we also consider a $240/8/1$ Walker-Star constellation system for comparison. The inclination of the Walker-Delta constellation is 45 degree and that of Walker-Star constellation is 99.5 degree. The altitude of the Walker-Delta orbits is 500 km and that of Walker-Star constellation is 700 km. Fig. 5 illustrates the virtual space-ground integrated system constructed for our simulations, encompassing terrestrial devices, different LEO satellite constellations, and GEO satellites. Simulation system with a $240/8/1$ Walker-Star constellation can be extended from Fig. 5(b).
@@
(a) Simulation system with a 80/4/1 Walker-Delta constellation.
(b) Simulation system with a 80/4/1 Walker-Star constellation.
Figure 5: Simulation system with different LEO constellations.
IV-A2 Learning Tasks
The batch-size is set to be 32. The learning rate is set to be $1\times 10^{-3}$. The number of communication rounds $T$ and the number of local update $E$ are set to be $300$ and $5$, respectively. Three metrics are used for analyzing the learning performance, which are the loss, the coefficient of determinations ($R^{2}$), and the mean absolute error (MAE). We compare the proposed hierarchical learning with geographical located devices with the centralized training case, where all the data are collected and training on the cloud server. We run all the experiments on CUDA version 11.7, PyTorch version 2.0.0, and Python version 3.9. The terrestrial webcams are also marked in Fig. 5.
(a) Training loss versus number of communication rounds.
(b) $R^{2}$ versus number of communication rounds.
(c) Testing MAE versus number of communication rounds.
(a) Training loss in a 80/4/1 Walker-Delta constellation.
(b) Training loss in a 80/4/1 Walker-Star constellation.
Figure 7: Learning performance versus system energy consumption.
@@ IV-A4 Experimental Platforms
We use the STK version 11.6 tool to model the space-ground integrated networks. Specifically, the $\mathrm{Chain}$ and $\mathrm{Constellation}$ objects are used for modeling both the ground-space connection and space topology. Specifically, $\mathrm{Chain}$ objects that connects ground devices and the LEO constellations, i.e., the $\mathrm{Constellation}$ objects, are utilized to determine the terminal satellites receiving local models. Matlab-STK connectors is utilized to extract range information between LEO satellites for space topology construction. The scenario time period is set to be from 25 Jun 2024 02:00:00.000 UTCG to 26 Jun 2024 02:00:00.000 UTCG.
@@ IV-B1 Learning Performance versus Communication Rounds
The training loss, test $R^{2}$, and MAE versus communication rounds of the environmental parameter retrieval are shown in Fig. 6. As observed in Fig. 6, our proposed hierarchical learning architecture progressively approaches the optimal performance of centralized training as the number of communication rounds increases. The performance gap is attributed to the non-independent and identically distributed (non-IID) nature of the data generated at terrestrial devices in different geographic locations. Given that the hierarchical learning architecture does not require uploading end data to the cloud server, it effectively protects image privacy for the environmental monitoring applications.
@@ IV-B2 Overall Results Analysis
Table II illustrates the overall simulation results for the TAEER problem. The performance metrics are average energy consumption per time slot (J) and average outage probability for each ISL (%) after $T=300$ communication rounds. $\rho=1$ represents the case of error-free ISL and $\rho=0.999$ represents the case for solving $\mathscr{P}_{5}$, which takes outage probability into account. As clearly illustrated in Table II, in the default settings, our proposed TAEER algorithm consistently demonstrates the lowest energy consumption and outage probability, followed by the D-Merge method, with the Orbit-Greedy method exhibiting the highest energy consumption and outage probability. This is because the Orbit-Greedy method only considers intra-orbit ISLs, resulting in multiple high-energy consumption links transmitting to GEO satellites within the network. By solving the topology-aware DST problem, TAEER algorithm and D-Merge achieve better overall performance. For instance, in the simulation system with Walker-Delta constellation, TAEER demonstrates a significant energy reduction, up to 3.64% and 67.50%, and an outage probability reduction, up to 11.97% and 12.75% compared with D-Merge and Orbit-Greedy, respectively. We can conclude that the heuristic solution, i.e., TAEER, has the best performance compared to the simple approximation, i.e., D-Merge, and the Orbit-Greedy scheme.
Besides, under default settings, each ISL transmission latency is on the millisecond level, while the length of each time frame is $\Delta\tau=10$ seconds. Therefore, if an ISL transmission happens to experience an transmission outage within a time frame, it can be quickly retransmitted without affecting the aggregation results. Additionally, the ISL capacity is at the Gbps level, capable of accommodating the transmission of larger-scale models, and it is possible to complete the transmission of an entire model data packet within a single time frame.
From $\rho=1$ to $\rho=0.999$ in Table II, we can see an increase in the average energy consumption. This is caused by the extra energy consumed by retransmission protocols in LTP when a transmission outage occurs at an ISL.
$\rho=1$ $\rho=0.999$
Average energy consumption per time slot (J) 80/4/1 Walker-Delta 180.68 187.50 555.73 182.81 196.93 556.38
80/4/1 Walker-Star 202.38 208.42 573.90 207.84 214.48 575.79
240/8/1 Walker-Star 151.20 153.87 552.79 151.31 154.14 553.90
Average outage probability for each ISL (%) 80/4/1 Walker-Delta - - - 3.97 4.51 4.55
80/4/1 Walker-Star - - - 5.36 6.22 6.35
240/8/1 Walker-Star - - - 1.35 1.39 1.47
IV-B3 Learning Performance versus System Energy Consumption
The training loss versus energy consumption in different LEO networks of the environmental parameters retrieval task is illustrated in Fig. 7. We can conclude that the proposed solutions, compared to Orbit-Greedy, which does not consider inter-orbit ISLs, significantly reduce the energy consumption in the LEO satellite network required to achieve the same learning accuracy across two different constellation configurations.

### v2

@@ III-B Problem Formulation
where $w_{i,j}(m)$ can be regarded as the weight of edge $(i,j)$ in $\mathcal{E}^{\prime}(m)$. The objective is to find a minimum directed sub-branching covering all the terminals with the lowest energy consumption. The constraint is that the edges of the sub-branching must be part of the directed graph $\bm{G}(m)$ constructed by the current network topology. Essentially, it involves identifying an inter-satellite routing path at minimum energy cost that takes the form of a directed sub-branching.
@@
to minimize the energy consumption in each time frame. Here $\mathcal{B}(m,u)$ denotes the directed sub-branching of each time frame. The constraint is that the edges of the directed sub-branching must be part of the subgraph $\bm{G}(m,u)$ constructed in each time frame. This means that the solution to $\mathscr{P}_{1}$ can be obtained by calculating the solution to each time frame by solving $\mathscr{P}_{2}$. It is worth noting that the local information on each satellite is assumed to be partitioned into $U$ parts, so that each part can be transmitted within the duration of a time frame. In this case, the energy consumption for ISLs in each time frame is given by
@@
Based on these assumptions, $\mathscr{P}_{2}$ is a static problem within each time frame. The solution set to $\mathscr{P}_{2}$ is not finite because, although the operation of LEO constellations exhibits periodicity, the Earth’s rotation causes the identity of terminals to change across communication rounds in the hierarchical learning procedures. The overall characterization of space topology is demonstrated in Fig. 4.
@@ III-C Proposed Topology-Aware Energy-Efficient Routing
    Shortest Path: Find the shortest distance path from each terminal node $i\in\mathcal{V}^{\prime}\setminus r$ to the root node $j=r$ using the Dijkstra algorithm [36], and record each predecessor node such that each terminal can trace back to the root node. This is because, compared to other similar algorithms, Dijkstra’s algorithm exhibits advantages in stability and computational efficiency for solving single-source shortest-path problems on graphs with non-negative edge weights. Construct a set $\mathrm{Path}=\{i\rightarrow r|i\in\mathcal{V}\setminus r\}$ consisting of node and edge information for all shortest paths.
@@
The detailed procedures of the proposed TAEER algorithm are presented in Algorithm 1, an simple example of the solution is illustrated in Fig. 5, and an example flowchart of Chu-Liu-Edmonds algorithm for solving the MSA problem is illustrated in Fig. 6. Based on the derived MSA $\mathcal{A}^{\ast}(m,u)$, a tree structure in such an LEO satellite network enables transmission-and-aggregation from each leaf node to the root node, ultimately completing the global model aggregation task.
@@ III-D Extension to Unpredictable Topology
When solving $\mathscr{P}_{1}$ and $\mathscr{P}_{2}$, we assume that the transmission over each hop of the laser ISL is perfect. However, in practical systems, a pointing, acquisition, and tracking (PAT) system is required to maintain precise link alignment to ensure the LOS characteristic. Satellite positioning inaccuracies, environmental disturbances, and mechanical vibrations are the main contributing factors to the misalignment of the telescope beam [42, 53]. Misalignment of the telescope beam can lead to pointing errors [42, 43], which in turn may cause transmission outages. This introduces uncertainties that we aim to address. We consider transmission failure caused by outage in the laser ISL. To characterize the outage probability of the laser ISL, it can be modeled as the probability that the signal-to-noise ratio (SNR) falls below a certain threshold, which can be expressed as
@@
To compute the overall outage probability of an aggregation path, which contains multi-hop laser ISLs, we first compute its success probability as
$P_{\text{succ}}(\mathcal{B}(m,u))=\prod_{i,j\in\mathcal{V}^{\prime}}[1-p_{\text{out}}^{i,j}(m,u)].$ (36)
@@
$\log P_{\text{succ}}(\mathcal{B}(m,u))=\sum_{i,j\in\mathcal{V}^{\prime}}\log[1-p_{\text{out}}^{i,j}(m,u)].$ (37)
Therefore, minimizing the outage probability of the aggregation path in each time frame to address the uncertainty is equivalent to minimizing $-\log P_{\text{succ}}(\mathcal{B}(m,u))$, which yields the reformulated optimization problem:
$\mathscr{P}_{3}:~{}\operatorname{minimize}~{}$ $\sum_{i,j\in\mathcal{V}^{\prime}}\log\left[\frac{1}{1-p_{\text{out}}^{i,j}(m,u)}\right].$ (38)
@@
$\mathscr{P}_{4}:~{}\operatorname{minimize}~{}$ $\rho W_{u}(\mathcal{B}(m,u))-(1-\rho)\log P_{\text{succ}}(\mathcal{B}(m,u)).$ (39)
It is worth noting $\mathscr{P}_{4}$ can be directly solved by adjusting the weight of the directed graph to $w^{r}_{i,j}=\rho w_{i,j}+(1-\rho)\log\frac{1}{1-P_{\text{out}}^{i,j}}$ using Algorithm 1.
[…]
III-E Convergence Results
To analyze convergence performance of the proposed hierarchical learning algorithm, we have the following assumptions and theorem.
Assumption 1.
The objective function $f$ is lower bounded by a scalar $f^{\inf}$, i.e., $f(\bm{x})\geq f^{\inf}>-\infty$.
Assumption 2 ($L$-Smoothness).
The loss function of each device $i$ in cluster $j$, each cluster $j$, and all clusters, i.e., $f_{i,j}(\cdot)$, $f_{j}(\cdot)$, and $f(\cdot)$ are all $L$-smooth. For all $\bm{x}$ and $\bm{y}$, we have
$\|\nabla f_{i,j}(\bm{x})-\nabla f_{i,j}(\bm{y})\|_{2}\leq L\|\bm{x}-\bm{y}\|,$
and
$f(\bm{y})\leq f(\bm{x})+\left\langle\nabla f(\bm{x}),\bm{y-x}\right\rangle+\frac{L}{2}\left\|\bm{y-x}\right\|_{2}^{2}.$
Assumption 3 (Unbiased Gradient and Bounded Variance).
Each device has an unbiased stochastic gradient estimation with bounded variance, i.e.,
$\mathbb{E}[\tilde{\nabla}f_{i,j}(\bm{x})]=\nabla f_{i,j}(\bm{x}),$
$\mathbb{E}\left[\|\tilde{\nabla}f_{i,j}(\bm{x})-\nabla f_{i,j}(\bm{x})\|_{2}^{2}\right]\leq\sigma^{2}.$
Assumption 4 (Bounded Gradient Dissimilarity).
There exists constants $\alpha\geq 1$ and $\beta\geq 0$ such that
$\frac{1}{J}\sum_{n=1}^{J}\|\nabla f_{n}(\bm{x})\|_{2}^{2}\leq\beta+\alpha\|\nabla f(\bm{x})\|_{2}^{2}.$ (40)
Assumptions 1, 2, and 3 are commonly used in stochastic optimization for the objective function [58]. Assumption 4 is adopted to characterize the non-IID extent of the local data distribution [59]. Similar assumptions have also been made in the literature [60, 61] for the convergence analysis under the non-i.i.d. setting.
Theorem 1 (Error Bound of Hierarchical Learning Under Nonconvexity).
Let Assumptions 1, 2, 3, and 4 hold. When the learning rate is bounded as
$\eta=\frac{1}{L}\sqrt{\frac{J}{ET}}\leq\min\left\{\frac{1}{2LE},\frac{1}{L\sqrt{2E(E-1)(2\alpha+1)}}\right\},$
the minimal gradient norm of the global loss function after $T$ communication rounds is bounded by
$\min_{t\in[T]}\mathbb{E}\left[\|\nabla f(\bm{x}^{t})\|_{2}^{2}\right]\leq$ $\mathcal{O}\left(\frac{1+\bar{\sigma}^{2}}{\sqrt{JET}}\right)$
$+$ $\mathcal{O}\left(\frac{J(E-1)(\sigma^{2}+2E\beta)}{ET}\right),$ (41)
where $\mathcal{O}(\cdot)$ omits the constants and $\bar{\sigma}^{2}=J\sum_{n=1}^{J}\sigma^{2}$.
Proof.
It is worth noting that the aggregation is based on the average of all devices’ local information. Based on the updating rules, i.e.,
$\bm{x}^{t+1}:=\bm{x}^{t}-\eta\sum_{n=1}^{J}\sum_{e=0}^{E-1}\lambda_{n}\tilde{\nabla}f_{n}(\bm{x}_{n}^{t,e}),$
and the authors previous results in [57, Supplementary Material], the average norm of global gradients after $T$ communication rounds can be bounded by
$\frac{1}{T}\sum_{t=0}^{T}\mathbb{E}\left[\|\nabla f(\bm{x}^{t})\|_{2}^{2}\right]$
$\leq$ $\frac{4[f(\bm{x}^{0})-f^{\inf}]}{\eta ET}+4\eta L\sigma^{2}\sum_{n=1}^{J}\lambda_{n}^{2}+3\eta^{2}L^{2}\sigma^{2}(E-1)$
$+$ $6\eta^{2}L^{2}\beta E(E-1).$ (42)
[…]
From Theorem 1, we can conclude that the convergence rate of the hierarchical learning algorithm in this paper is decided by $\mathcal{O}\left(\frac{1+\bar{\sigma}^{2}}{\sqrt{JET}}\right)$, which presents a linear speedup caused by $E$ rounds of local updates. An additional error is caused by data heterogeneity and multiple local updates, which means the proposed algorithm has an error gap compared with centralized training.
@@ IV-A1 Network Configurations
To verify the performance of the proposed algorithms, we consider a $80/4/1$ Walker-Star constellation system and a $80/4/1$ Walker-Delta constellation system, which consists of $N=4$ orbit planes and $S_{n}=20$ satellites evenly distributed in each orbit for default settings. Three GEO satellites are evenly distributed in the Equatorial plane orbit. In order to characterize the impact of the network size, we also consider a $240/8/1$ Walker-Star constellation system for comparison. We also consider a $800/20/1$ large Walker-Star constellation system to simulate the real-world network scale. The inclination of the Walker-Delta constellation is 45 degree and that of Walker-Star constellation is 99.5 degree. The altitude of the Walker-Delta orbits is 500 km and that of Walker-Star constellation is 700 km. Fig. 7 illustrates the virtual space-ground integrated system constructed for our simulations, encompassing terrestrial devices, different LEO satellite constellations, and GEO satellites. Simulation system with $240/8/1$ and $800/20/1$ Walker-Star constellation can be extended from Fig. 7(b).
(a) Simulation system with a 80/4/1 Walker-Delta constellation.
(b) Simulation system with a 80/4/1 Walker-Star constellation.
Figure 7: Simulation system with different LEO constellations.
(a) Training loss versus number of communication rounds.
(b) $R^{2}$ versus number of communication rounds.
(c) Testing MAE versus number of communication rounds.
IV-A2 Learning Tasks
The batch-size is set to be 32. The learning rate is set to be $1\times 10^{-3}$. The number of communication rounds $T$ and the number of local update $E$ are set to be $300$ and $5$, respectively. Three metrics are used for analyzing the learning performance, which are the loss, the coefficient of determinations ($R^{2}$), and the mean absolute error (MAE). We compare the proposed hierarchical learning $E=\{1,5,10\}$ with geographical located devices with the centralized training case, where all the data are collected and training on the cloud server to verify the convergence results. We run all the experiments on CUDA version 11.7, PyTorch version 2.0.0, and Python version 3.9. The terrestrial webcams are also marked in Fig. 7.
@@ IV-A4 Experimental Platforms
We use the STK version 11.6 tool to model the SGINs. Specifically, the $\mathrm{Chain}$ and $\mathrm{Constellation}$ objects are used for modeling both the ground-space connection and space topology. Specifically, $\mathrm{Chain}$ objects that connects ground devices and the LEO constellations, i.e., the $\mathrm{Constellation}$ objects, are utilized to determine the terminal satellites receiving local models. Matlab-STK connectors is utilized to extract range information between LEO satellites for space topology construction. The scenario time period is set to be from 25 Jun 2024 02:00:00.000 UTCG to 26 Jun 2024 02:00:00.000 UTCG.
The parameter settings in this paper are summarized in Table I, where the selection of parameters is primarily based on [38, 43, 44, 42, 46]. The simulation code of this paper under a Walker Star constellation available at https://github.com/Jingyang-Zhu/HierLC as a demonstration.
@@ IV-B1 Learning Performance versus Communication Rounds
The training loss, test $R^{2}$, and MAE versus communication rounds of the environmental parameter retrieval are shown in Fig. 8. As observed in Fig. 8, our proposed hierarchical learning architecture progressively approaches the optimal performance of centralized training as the number of communication rounds increases. The performance gap is attributed to the non-independent and identically distributed (non-IID) nature of the data generated at terrestrial devices in different geographic locations. Given that the hierarchical learning architecture does not require uploading end data to the cloud server, it effectively protects image privacy for the environmental monitoring applications. In addition, to demonstrate the impact of the number of local updates on the hierarchical learning architecture, as shown in Fig. 8, the learning performance and convergence speed under $E=10$ outperform those with $E=5$ and $E=1$, which aligns with the conclusions of our convergence analysis.
(a) Training loss in a 80/4/1 Walker-Delta constellation.
(b) Training loss in a 80/4/1 Walker-Star constellation.
Figure 9: Learning performance versus system energy consumption.
$\rho=1$ $\rho=0.1$
Average energy consumption per time slot (J) 80/4/1 Walker-Delta 180.68 187.50 555.73 183.90 221.41 556.38
80/4/1 Walker-Star 202.38 208.42 573.90 211.49 276.25 575.79
240/8/1 Walker-Star 151.20 153.87 552.79 151.97 161.24 552.87
800/20/1 Walker-Star 160.21 169.86 542.78 160.51 182.03 542.78
Average outage probability for each ISL (%) 80/4/1 Walker-Delta - - - 6.71 9.70 5.37
80/4/1 Walker-Star - - - 8.97 15.3 7.50
240/8/1 Walker-Star - - - 1.56 2.06 1.47
800/20/1 Walker-Star - - - 0.75 1.07 0.92
@@ IV-B2 Overall Results Analysis
Table II illustrates the overall simulation results for the TAEER problem. The performance metrics are average energy consumption per time slot (J) and average outage probability for each ISL (%) after $T=300$ communication rounds. $\rho=1$ represents the case of error-free ISL and $\rho=0.1$ represents the case for solving $\mathscr{P}_{4}$, which takes outage probability into account. As clearly demonstrated in Table II, under the default settings, our proposed TAEER algorithm consistently achieves the lowest energy consumption, followed by the D-Merge method, while the Orbit-Greedy method exhibits the highest energy consumption. With respect to outage probability, Orbit-Greedy yields the lowest value, followed by TAEER, with D-Merge showing the highest. This disparity arises because the Orbit-Greedy method exclusively utilizes stable intra-orbit ISLs, leading to multiple high-energy consumption links transmitting to GEO satellites within the network. In contrast, both TAEER and D-Merge account for unstable inter-orbit ISLs. By solving the topology-aware DST problem, TAEER algorithm and D-Merge achieve better overall performance. For instance, in the simulation system with Walker-Delta constellation, TAEER demonstrates a significant energy reduction, up to 3.64% and 67.50% compared with D-Merge and Orbit-Greedy, and a 30.82% outage probability reduction compared with D-Merge. We can conclude that the heuristic solution, i.e., TAEER, has the best performance compared to the simple approximation, i.e., D-Merge, and the Orbit-Greedy scheme.
Besides, under default settings, each ISL transmission latency is on the millisecond level, while the length of each time frame is $\Delta\tau=10$ seconds. Therefore, if an ISL transmission happens to experience an transmission outage within a time frame, it can be quickly retransmitted without affecting the aggregation results. We calculate the time required to solve $\mathscr{P}_{2}$ with $800/20/1$ configuration, the convergence time of the proposed TAEER algorithm is 0.027 seconds, while that of D-Merge is 0.003 seconds. Both are significantly shorter than $\Delta\tau=10$ seconds. Therefore, our proposed algorithms can meet the real-time requirements of satellite network applications. Additionally, the ISL capacity is at the Gbps level, capable of accommodating the transmission of larger-scale models, and it is possible to complete the transmission of an entire model data packet within a single time frame.
From $\rho=1$ to $\rho=0.1$ in Table II, we can see an increase in the average energy consumption. This is caused by the extra energy consumed by retransmission protocols in LTP when a transmission outage occurs at an ISL.
@@
IV-B3 Learning Performance versus System Energy Consumption
The training loss versus energy consumption in different LEO networks of the environmental parameters retrieval task is illustrated in Fig. 9. To analyze the performance of the proposed TAEER and the D-Merge, we first conclude that D-Merge is a well-established approximation algorithm proposed for solving the DST problem, which has the similar performance as the proposed TAEER. While this approach has low complexity, it may lead to suboptimal performance due to insufficient path sharing. TAEER method constructs subgraphs and fully leverages the benefits of global path sharing and relay nodes to enhance the performance. We shall provide further comparisons between these two schemes in the following subsections. We can also conclude that the proposed solutions, compared to Orbit-Greedy, which does not consider inter-orbit ISLs, significantly reduce the energy consumption in the LEO satellite network required to achieve the same learning accuracy across two different constellation configurations.